# Augmenting LLMs for General Time Series Understanding and Prediction

### v1

@@ 3.2 TsLLM Architecture
Having established how we encode time series into dense representations, we now describe how these representations are integrated with a pretrained language model to create a unified multimodal system. Our architectural design faces a fundamental challenge: how to combine time series understanding with the rich language capabilities of modern LLMs without degrading either modality’s performance.
Foundation Model Integration
Rather than training a multimodal model from scratch, we build upon existing pretrained LLMs for both practical and technical reasons. First, the scarcity of high-quality time series data on the internet—our analysis shows less than 0.1% of web text contains meaningful temporal sequences—makes joint pretraining economically infeasible. Second, pretrained LLMs already possess sophisticated reasoning capabilities, world knowledge, and instruction-following abilities that would be difficult to replicate. Our architecture therefore, focuses on augmenting these capabilities rather than replacing them.
@@
We designed TsLLM to be model-agnostic, compatible with any decoder-only transformer architecture. In our experiments (Sect. 4), we demonstrate strong results with Qwen2.5 (Wang et al., 2024), Llama 3.2 (Grattafiori et al., 2024), and SmolLM2 (Allal et al., 2024), showing that our approach generalizes across different model families and scales, and that TsLLM can leverage future improvements in base LLM capabilities to enhance its performance.
Adapter Layers
The adapter layers serve as the critical bridge between our time series representations and the LLM’s embedding space. These components must solve a challenging alignment problem: mapping from the encoder’s learned representations (optimized for reconstruction) to the LLM’s token embeddings (optimized for next-token prediction). Our adapter architecture employs a two-layer MLP with residual connections:
$\mathbf{h}_{\text{adapted}}=\mathbf{z}+\text{MLP}_{2}(\text{GELU}(\text{MLP}_{1}(\mathbf{z})))$
where $\mathbf{z}$ is the encoded time series representation. The residual connection preserves the original temporal information while the MLP learns the necessary transformations for LLM compatibility. During training, we keep the encoder frozen and only update the adapter weights, preventing catastrophic forgetting of the encoder’s learned representations. We experimented with more sophisticated architectures, including cross-attention mechanisms (where time series representations query LLM embeddings) and Perceiver-based adapters (uses learned queries to extract information). However, ablation studies showed that simple MLPs performed comparably with 3x fewer parameters and 2x faster inference. This efficiency is crucial for practical deployment, especially for long time series sequences.
Scale-Shape Decomposition
@@
Time series in real-world applications span vastly different scales, e.g., heartbeats measured in millivolts, stock prices in thousands of dollars, and particle counts in millions. This variability presents unique challenges for multimodal models. Unlike traditional time series models that can normalize inputs and rescale outputs using known statistics, TsLLM must: (1) understand scale-dependent semantics (e.g., a heart rate of 200 bpm indicates tachycardia), and (2) generate time series at appropriate scales from text-only prompts (e.g., “Generate normal ECG signals”). Our solution decouples each time series into shape (normalized patterns) and scale (magnitude information) components. While the encoder processes normalized sequences $\hat{\mathbf{x}}=(\mathbf{x}-\mu_{\mathbf{x}})/\sigma_{\mathbf{x}}$ to extract shape features, we separately encode the normalization parameters $\mu_{\mathbf{x}}$ and $\sigma_{\mathbf{x}}$ to preserve scale information.
Model Fusion
Integrating time series representations into the LLM requires careful handling of the token vocabulary and embedding spaces. We extend the LLM’s vocabulary with a special “<|ts|>” token that acts as a placeholder for time series patches. During preprocessing, we identify time series segments in the input and replace each patch with a “<|ts|>” token, maintaining the sequential structure of the data. At the embedding layer, we implement a routing mechanism that replaces “<|ts|>” token embeddings with their corresponding time series representations:
$\mathbf{e}_{i}=\begin{cases}\text{TokenEmbed}(t_{i})&\text{if }t_{i}\neq\texttt{<|ts|>}\\ \mathbf{h}_{\text{adapted},j}+\mathbf{s}_{j}&\text{if }t_{i}=\texttt{<|ts|>}\end{cases}$
where $\mathbf{h}_{\text{adapted},j}$ is the $j$-th adapted time series representation and $\mathbf{s}_{j}$ is its corresponding scale embedding. This additive combination allows the model to modulate time series patterns based on their scale context. For generation, we implement bidirectional conversion: the LLM predicts “<|ts|>” tokens where time series should be generated, and we extract the corresponding hidden states for decoding. This design enables truly interleaved generation—the model can freely mix text and time series in its outputs based on the task requirements.
@@ 3.3 Training TsLLM
Figure 2: Illustrative examples of our unified framework where time series data is integrated with the text in both the inputs and outputs. This figure visualizes these embedded time series segments as plots for illustrative clarity. Examples demonstrate versatility across tasks like: (Top) contextual forecasting (e.g., ECG signals), (Middle) classification (e.g., sensor data), and (Bottom) report generation (e.g., COVID mobility).
Unified Task Formulation
A key feature of TsLLM is reformulating all time series tasks as autoregressive token generation problems. This unification enables knowledge transfer across tasks and simplifies the training process. Forecasting becomes natural continuation: given historical data as time series tokens, the model generates future time series tokens. The autoregressive nature means every time series in our training data implicitly teaches forecasting. Classification transforms into constrained text generation. Given a time series and the prompt “What type of activity is this?”, the model generates class labels as text. We use constrained decoding during inference to ensure outputs match valid class names. Anomaly detection employs two complementary approaches: (1) direct detection where the model outputs anomaly locations as text (eg. “Anomalies detected at timestamps 145-162, 203-210”), and (2) reconstruction-based detection where the model generates expected normal behavior, with deviations indicating anomalies.
Other tasks including segmentation, point-level classification, and imputation can similarly be formulated as text generation tasks. This unified formulation is powerful but introduces a challenge: dataset-specific semantics. A “normal” pattern in one domain may be anomalous in another. We address this through few-shot prompting, including domain examples in the context to provide task-specific guidance without fine-tuning.
Multi-Objective Training
Our training objective balances three components: $\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{\text{CE}}+\lambda_{2}\mathcal{L}_{\text{MSE}}+\lambda_{3}\mathcal{L}_{\text{scale}}.$ The cross-entropy loss $\mathcal{L}_{\text{CE}}$ applies to text tokens, maintaining language modeling capabilities. The MSE loss $\mathcal{L}_{\text{MSE}}$ supervises time series reconstruction through the decoder. The scale loss $\mathcal{L}_{\text{scale}}$ ensures accurate recovery of normalization parameters. We use adaptive weighting with $\lambda_{1}=1.0$, $\lambda_{2}=0.5$, and $\lambda_{3}=0.1$, determined through extensive hyperparameter search to balance modalities without one dominating.
@@
The training process for TsLLM employs a carefully orchestrated three-stage approach, with each stage serving a specific purpose in building the model’s capabilities. Stage 1: Alignment (1B tokens) – We begin by training only the adapter and scale encoding layers while keeping both the encoder-decoder and LLM frozen. This stage uses raw time series sequences and simple synthetic tasks (estimating the mean, translating to and from text-based time series encodings, etc.) to establish the connection between modalities. By limiting the trainable parameters, we ensure stable alignment without disrupting pretrained representations. Stage 2: Continued Pretraining (8B tokens) – With alignment established, we unfreeze all parameters for comprehensive multimodal training. This stage uses our full training dataset, with conversational prompt-response data undersampled and reformatted as unstructured sequences. Stage 3: Supervised Fine-Tuning (20B tokens) – The final stage shifts to a conversational format, training on instruction-response pairs with loss computed only on responses. This stage transforms TsLLM from a base model for next-token prediction into an instruction-following model that can solve time series analysis tasks specified in natural language. This staged approach proved crucial for stability as early experiments with all parameters unfrozen from the start tended to have loss spikes, and performance plateaued early.
@@ 3.4 Data Curation and Synthesis
Training TsLLM to bridge time series and language modalities requires comprehensive data coverage across three categories: pure time series for encoder pretraining, paired time series with text for multimodal alignment, and synthetic tasks incorporating time series with natural language instructions and context for capability development. Our training corpus comprises 25 billion tokens across 2 million examples, carefully curated to address the scarcity of high-quality time series-text pairs while ensuring broad domain coverage. The data collection process addresses a fundamental challenge: while web-scraped text corpora contain billions of documents, less than 0.1% include meaningful temporal sequences, necessitating alternative strategies for collecting suitable data.
Existing Datasets
We leverage established benchmarks to provide foundational capabilities across diverse time series tasks. For large-scale pretraining, we utilize GIFT-Eval (Aksu et al., 2024) with 230 billion time series points spanning 88 domains including healthcare, finance, energy, and environmental monitoring. This scale ensures our encoder-decoder learns robust temporal representations across different sampling frequencies (from seconds to years) and sequence characteristics. Classification capabilities are developed using the UCR (Dau et al., 2018) and UEA (Bagnall et al., 2018) archives, together covering 158 datasets ranging from physiological signals like electrocardiograms to industrial sensor readings and human activity recognition.
For paired time series-text data, we incorporate specialized datasets that naturally combine both modalities. ECG-QA (Oh et al., 2023) provides 200,000 cardiac signal question-answering pairs spanning diagnostic queries, rhythm identification, and clinical interpretation tasks. The TextTime Corpus (Kim et al., 2024) contributes weather measurements paired with meteorological discussions and clinical notes synchronized with patient vital signs from intensive care units. These existing paired datasets, while valuable, represent only a small fraction of our training data, highlighting the need for augmentation strategies.
Dataset Augmentation
To expand beyond limited existing paired datasets, we develop novel augmentation strategies that create meaningful text-time series associations while preserving signal fidelity. Our financial dataset pairs 5 years of S&P 500 price histories with 62,000 contemporaneous news articles from financial news sources, temporally aligned to enable contextual forecasting where predictions must integrate both numerical trends and market events. Each example combines a 32-day price history with up to 4 full articles or 16 headlines, ensuring textual context precedes prediction windows.
For medical applications, we augment the PTB-XL ECG dataset (Wagner et al., 2022) containing 21,799 clinical recordings to create three specialized variants through controlled text generation. The contextual forecasting variant pairs each ECG with multiple clinical reports using templates ranging from basic summaries to detailed interpretative analyses. The classification variant combines signals with demographic narratives and symptom descriptions while excluding pre-computed features, forcing reliance on raw signal patterns. The report generation variant produces structured findings, narrative clinical stories, and differential diagnoses from ECG patterns. This augmentation strategy transforms a single dataset into 65,000 multimodal training examples covering diverse analytical perspectives.
Synthetic Data Generation
The scarcity of naturally occurring paired time series-text data necessitates large-scale synthesis to achieve comprehensive task coverage. We generate 2 million question-answer pairs (10 billion tokens) through systematic combination of time series samples with templated analytical tasks. Our synthesis pipeline samples time series from diverse domains, constructs task-specific prompts spanning trend analysis, anomaly detection, periodicity identification, statistical reasoning, and causal interpretation, then generates responses that provide detailed analytical narratives.
Beyond question-answering, we create foundational tasks that establish core capabilities: bidirectional conversion between numerical arrays and textual representations (testing precise value preservation), statistical computations ranging from simple aggregations to complex distributional analyses, and multi-scale temporal reasoning from local patterns to long-term trends. This synthetic data follows recent findings (Bansal et al., 2024) suggesting that larger quantities of diverse examples, even with occasional inaccuracies, provide more robust training signals than smaller curated datasets. Our synthesis process ensures coverage of edge cases, rare patterns, and complex reasoning chains that might be underrepresented in natural data.
Data Formatting and Task Unification
We unify diverse time series tasks through consistent prompt-response formatting that enables seamless interleaving of text and time series. As illustrated in Figure 2, each task follows a structured template combining natural language instructions, optional context, time series signals (represented as <|ts|> tokens), and expected outputs. For contextual forecasting, prompts include historical signals, metadata, and relevant textual context, with responses generating future time series continuations. Classification tasks provide dataset descriptions and few-shot examples, enabling in-context learning without task-specific fine-tuning. Report generation prompts combine time series with metadata, producing detailed textual analyses that describe patterns, identify anomalies, and provide domain-specific interpretations. This unified formatting allows TsLLM to treat all tasks as variations of autoregressive generation, leveraging its language modeling capabilities while maintaining precise numerical fidelity. Complete details on dataset composition, synthesis procedures, and formatting specifications are provided in Appendix C.
@@ 4 Results
To demonstrate TsLLM’s unique capabilities at the intersection of time series analysis and natural language understanding, we evaluate it across diverse tasks ranging from those requiring language integration (question-answering, contextual forecasting) to traditional time series benchmarks (classification, anomaly detection). This evaluation strategy reflects our aims: while we show competitive performance on traditional tasks, our primary focus is demonstrating capabilities that existing approaches cannot provide. We prioritize evaluation on time series question-answering, where TsLLM’s ability to combine temporal reasoning with domain knowledge is most evident. We also assess performance on traditional benchmarks to establish that augmenting an LLM with time series perception maintains competitive numerical analysis capabilities. The datasets span diverse domains including healthcare, finance, and industrial applications, with detailed anomaly detection results presented in Table 7 in the Appendix.
@@ 4.1 Evaluations
To evaluate the time series reasoning and question-answering capabilities of TsLLM, we assess its performance using the electrocardiogram question-answering (ECG-QA) benchmark (Oh et al., 2023) comprised of ECG recordings, questions, and answers across several question types: yes/no verification (single-verify), binary choice selection (single-choose), and open-ended queries (single-query). We compared TsLLM performance (zero-shot and few-shot) based on exact match accuracy to top-performing specialized models trained on the ECG-QA dataset, such as M3AE (Chen et al., 2022), MedViLL (Oh et al., 2023), SE-WRN + GPT-4 (Oh et al., 2023), Q-Heart (Pham et al., 2025), and ECG-LM (Yang et al., 2025). All these models are explicitly trained for ECG-QA, whereas TsLLM is exposed to ECG-QA during training, but only a small fraction (4%) of its training tokens originated from this dataset. The performance of TsLLM is further enhanced through in-context few-shot learning using just four examples from the same question type. As Table 1 illustrates, despite the lack of fine-tuning, the models perform well in the category, and TsLLM (few-shot) achieves the best or second-best performance across all queries. While the table focuses on a subset of the question types from this dataset, TsLLM can be applied to other question types with the proper prompt format, demonstrating the ability of the models to handle complex tasks that require expertise.
Table 1: QA accuracy comparison on ECG-QA. The best performance is in bold and the second best is underlined. TsLLM, particularly with few-shot prompting, achieves high performance despite not being fine-tuned for ECG-QA.
Question M3AE MedViLL SE-WRN Q-Heart ECG-LM TsLLM TsLLM
+ GPT-4 (few-shot)
S-Verify 74.6 73.9 71.0 90.9 75.8 77.1 85.3
S-Choose 57.1 54.1 48.1 62.6 57.4 58.9 60.8
S-Query 41.0 40.4 35.7 32.9 39.9 48.2 49.0
Average 57.6 56.1 51.6 61.4 57.7 61.4 65.0
We also evaluate TsLLM on TimeSeriesExam (Cai et al., 2024), a multiple-choice question answering benchmark that tests LLMs’ time series understanding across diverse settings. TsLLM performs far better than even frontier models many times larger than it, demonstrating the effectiveness of our encoding and training approach.
Table 2: Accuracy on TimeSeriesExam.
@@
Classification
For classification tasks, we evaluate our model on the UCR-TSC dataset. Performance is assessed using the mean and median accuracy across sub-datasets, ensuring robustness. Results can be found in Table 3.

### v2

@@ 3.2 TsLLM Architecture
Having established how we encode time series into dense representations, we now describe how these representations are integrated with a pretrained LLM to create a unified multimodal model. Our architectural design faces a fundamental challenge: how to combine time series and natural language modeling without degrading either modality’s performance.
@@
LLM Integration   Rather than training a multimodal model from scratch, which is computationally infeasible, we build upon existing pretrained LLMs. We designed TsLLM to be model-agnostic, compatible with any decoder-only transformer architecture. In our experiments (Sect. 5), we demonstrate strong results with Qwen2.5 [26], Llama 3.2 [27], and SmolLM2 [28], showing that our approach generalizes across different model families and scales.
Adapter Layers   The adapter layers serve as the critical bridge between our time series representations and the LLM’s embedding space. These components must solve a challenging alignment problem: mapping from the encoder’s learned representations (optimized for reconstruction) to the LLM’s token embeddings (optimized for next-token prediction). Our encoder adapter architecture utilizes cross-attention between the patch encodings and a set of learned embeddings. Formally, we compute $y=\text{softmax}((xq)k^{T})v$ where $x\in\mathbb{R}^{n\times d_{t}}$ is the input sequence of $n$ tokens, $q\in\mathbb{R}^{d_{t}\times d_{L}}$ is a linear mapping to the embedding space of the LLM, and $k,v\in\mathbb{R}^{z\times d_{L}}$ are the $z$ learnable key-value embedding pairs. Essentially, this maps each patch embedding of the input time series to a (dynamic) convex combination of learned value embeddings. The key and value embeddings are initialized to a sample of $z=1024$ embeddings from the pretrained LLM’s input embedding layer. If frozen, this would mean the time series tokens are mapped directly into the space of the LLM’s text embeddings, minimizing the domain gap, and the training just allows it to learn this mapping between domains better. The decoder adapter layer uses a more traditional MLP architecture.
@@
Scale-Shape Decomposition   Time series in real-world applications span vastly different scales, e.g., heartbeats measured in millivolts, stock prices in thousands of dollars, and particle counts in millions. This variability presents unique challenges for multimodal models. Unlike traditional time series models that can normalize inputs and rescale outputs using known statistics, TsLLM must: (1) understand scale-dependent semantics (e.g., a heart rate of 200 bpm indicates tachycardia), and (2) generate time series at appropriate scales from text-only prompts (e.g., predict stock prices given a news report). Our solution decouples each time series into shape (normalized patterns) and scale (magnitude information) components. While the encoder processes normalized sequences $\hat{\mathbf{x}}=(\mathbf{x}-\mu_{\mathbf{x}})/\sigma_{\mathbf{x}}$ to extract shape features, we separately encode the normalization parameters $\mu_{\mathbf{x}}$ and $\sigma_{\mathbf{x}}$ to preserve scale information. We apply a log transformation to compress the scale range while preserving relative differences. The transformed parameters are encoded through dedicated MLPs and combined with the shape representations via addition. During generation, the model predicts both shape and scale for each token. We apply inverse transformations to recover the original scale, enabling generation of appropriately-scaled time series.
Model Fusion   Integrating time series representations into the LLM requires careful handling of the token vocabulary and embedding spaces. We extend the LLM’s vocabulary with a special “<|ts|>” token that acts as a placeholder for time series patches. During preprocessing, we identify time series segments in the input and replace each patch with a “<|ts|>” token, then in the embedding step we replace “<|ts|>” token embeddings with their corresponding time series representations. The embedding step is then given by: $\mathbf{e}_{i}=\begin{cases}\text{embed}(t_{i})&\text{if }t_{i}\neq\texttt{<|ts|>}\\ \mathbf{h}_{\text{adapted},j}+\mathbf{s}_{j}&\text{if }t_{i}=\texttt{<|ts|>}\end{cases}$ where $\mathbf{h}_{\text{adapted},j}$ is the $j$-th adapted time series representation and $\mathbf{s}_{j}$ is its corresponding scale embedding. For generation, we implement the same process in reverse: TsLLM is trained to predict “<|ts|>” tokens where time series is present, and we extract the corresponding hidden states for decoding. This design enables truly interleaved generation, meaning the model can freely mix text and time series in its outputs based on the task requirements.
@@ 3.3 Training TsLLM
Figure 2: Illustrative example prompt-response pairs from the training dataset that demonstrate how TsLLM can interleave text and time series segments into a single sequence. The time series segments are plotted in this figure for clarity, but are represented as a sequence of embeddings in the model. Examples demonstrate versatility across tasks like: contextual forecasting (top), classification (middle), and report generation (bottom).
Unified Task Formulation   A key feature of TsLLM is reformulating all time series tasks as autoregressive token generation problems over sequences of mixed text and time series tokens. This unification enables knowledge transfer across tasks and eliminates the need for task-specific or dataset-specific heads, simplifying training and inference. Figure 2 provides example prompts and responses for several tasks to demonstrate how this is done. Question-answering, report generation, and other text generation tasks require no special handling. Forecasting is done through autoregressive time series patch generation and for inference, each decoded patch is appended to the input and fed back into the model to predict the next patch. Classification maps onto constrained text generation; the model generates class labels directly as text. We use constrained decoding during inference to ensure outputs match valid class names. Anomaly detection can employ two approaches: direct detection, where the model outputs anomaly locations as text (eg. “anomalies detected at timestamps 145-162, 203-210”), and reconstruction-based detection, where the model generates a time series of the expected normal behavior, with deviations indicating anomalies. Other tasks, including segmentation, point-level classification, and imputation, can similarly be formulated as time series generation tasks.
Multi-Objective Training   Our training objective balances three components: $\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{\text{CE}}+\lambda_{2}\mathcal{L}_{\text{MSE}}+\lambda_{3}\mathcal{L}_{\text{scale}}.$ The cross-entropy loss $\mathcal{L}_{\text{CE}}$ applies to text tokens, maintaining language modeling capabilities. The MSE loss $\mathcal{L}_{\text{MSE}}$ supervises time series reconstruction through the decoder. The scale loss $\mathcal{L}_{\text{scale}}$ ensures accurate recovery of normalization parameters.
@@
The training process for TsLLM employs a carefully orchestrated three-stage approach, with each stage serving a specific purpose in building the model’s capabilities. The staged approach proved crucial for stability as early experiments with all parameters unfrozen from the start tended to have loss spikes and performance plateaued early. Stage 1: Alignment (1B tokens) – We begin by training only the adapter and scale encoding layers while keeping both the encoder-decoder and LLM frozen. This stage uses raw time series sequences and simple synthetic tasks (estimating the mean, translating to and from text-based time series encodings, etc.) to establish the connection between modalities. By limiting the trainable parameters, we ensure stable alignment without disrupting pretrained representations. Stage 2: Continued Pretraining (9B tokens) – With alignment established, we unfreeze all parameters for comprehensive multimodal training. This stage samples from the full training dataset, with conversational prompt-response data undersampled and reformatted as unstructured sequences. Stage 3: Supervised Fine-Tuning (20B tokens) – The final stage shifts to a conversational format, training on instruction-response pairs with loss computed only on responses. This stage transforms TsLLM from a base model for next-token prediction into an instruction-following model that can solve time series analysis tasks specified in natural language.
@@ 3.4 Data Curation and Synthesis
Training TsLLM to bridge time series and language modalities requires comprehensive data coverage across three categories: pure time series for encoder pretraining, paired time series with text for multimodal alignment, and synthetic tasks incorporating time series with natural language instructions and context for capability development. Our training corpus comprises approximately 27 billion tokens across 94 million examples, carefully curated to address the scarcity of high-quality time series-text pairs while ensuring broad domain coverage. Extensive work went into collecting and synthesizing sufficient high-quality data for training. This broadly included collecting and processing existing datasets (many with only time series, or time series plus labels), augmenting these datasets by connecting the signals with relevant text, and converting them into prompt-response pairs. It also involved generating synthetic prompt-response pairs for simple tasks and general QA, which make up a large portion of the training tokens. An in-depth description of the complete data gathering process is provided in Appendix B.
4 Implementation Details
The patch encoder and decoder each employed a 16-layer MLP with SwiGLU, LayerNorm, embedding size 128, and patch size 32. The VAE uses a 12-dimensional latent space with KL divergence scaled by peak $\beta=1.5$, trained for 100M samples (batch size 2048) using AdamW (peak LR 5e-4) and MSE reconstruction loss. TsLLM was optimized using AdamW (peak LR 5e-5) with cosine scheduling, linear warmup, and gradient clipping at 1.0. Training used 30B tokens sampled from a 27B token dataset with max sequence length 4096, applying MSE loss for time series prediction and cross-entropy for language tokens. We use Qwen2.5-7B [26] as the base LLM. Training required approximately 800 GPU-hours on four NVIDIA H100s. The added components total 6.7M parameters (1M each for encoder/decoder and scale modules, 3.7M for encoder adapter, 1M for decoder adapter), which is small relative to the base LLM.
@@ 5 Results
To demonstrate TsLLM’s unique capabilities at the intersection of time series analysis and natural language understanding, we evaluate it primarily on time series tasks that require language integration (question-answering, contextual forecasting). We provide additional results on traditional time series tasks (classification, forecasting, anomaly detection) in Appendix A to establish that it can achieve competitive performance with specialized models on these types of tasks even if it does not outright beat them. This evaluation strategy reflects our aims: while we show competitive performance on traditional tasks, our primary focus is demonstrating capabilities that existing approaches cannot provide.
@@ 5.1 Evaluations
To evaluate the time series reasoning and question-answering capabilities of TsLLM, we assess its performance using the electrocardiogram question-answering (ECG-QA) benchmark [29] comprised of ECG recordings, questions, and answers across several question types: yes/no verification (S-verify), choose between options (S-choose), and open-ended queries (S-query). We compared TsLLM performance based on exact match accuracy to top-performing specialized models trained on the ECG-QA dataset, such as M3AE [30], Q-Heart [31], and ECG-LM [32]. All these models are explicitly trained for ECG-QA, whereas TsLLM is exposed to ECG-QA during training, but only a small fraction (<4%) of its training tokens originated from this dataset, and its design is not specialized for ECG-QA. We also compare it to several LLMs and two time series LLMs, none of which were trained on ECG-QA, but receive few-shot examples in context. As Table 1 illustrates, despite the lack of fine-tuning, the models perform well in the category, and TsLLM (few-shot) earns the best or the second-best performance across all queries. While the table focuses on a subset of the question types from this dataset, TsLLM can be applied to other question types with the proper prompt format, demonstrating the ability of the models to handle complex tasks that require expertise.
Table 1: Question-answering exact-match accuracy comparison on ECG-QA by question type. Models with the “[Uncaptioned image]” icon have not been trained at all on ECG-QA, but received few-shot examples as context.
Question Type
Method S-Verify S-Choose S-Query Average
M3AE 74.6 57.1 41.0 57.6
Q-Heart 90.9 60.3 32.9 61.4
ECG-LM 75.8 57.4 39.9 57.7
Gemini 2.5 Pro [Uncaptioned image] 55.7 30.4 17.6 30.9
o4-mini [Uncaptioned image] 50.0 47.8 26.5 45.9
ChatTime [Uncaptioned image] 50.1 30.9 15.5 30.5
ChatTs [Uncaptioned image] 54.4 38.2 19.1 37.2
TsLLM 85.3 60.8 49.0 65.0
We also evaluate TsLLM on TimeSeriesExam [33], a multiple-choice question answering benchmark that tests LLMs’ time series understanding across diverse settings. TsLLM performs far better than even frontier models many times larger than it, demonstrating the effectiveness of our encoding and training approach.
Table 2: Accuracy on TimeSeriesExam, a synthetic time series question-answering benchmark.
@@
Contextual Forecasting
We evaluate the ability of TsLLM to forecast time series using the Context Is Key (CiK) dataset [34], which contains 71 realistic forecasting tasks spanning 7 domains specifically designed to require understanding and integrating textual information for successful prediction. TsLLM outperforms both traditional ML methods (which cannot use the context) as well as frontier LLMs (which can use the context, but have poor time series understanding) by a significant margin, demonstrating that it is able to bridge the gap between numerical signals and text-based context far better than existing methods.
Table 3: Contextual forecasting performance on the Context Is Key benchmark.
Model SMAPE (%) MASE (%)
Linear Regression 75.4 101.5
XGBoost 76.8 80.2
ARIMA 90.7 134.4
Qwen 2.5 7B 92.6 139.2
Gemini 2.5 Pro 90.8 98.1
o4-mini 72.6 70.5
ChatTime 70.1 71.4
ChatTs 68.4 69.2
TsLLM 64.5 64.7
5.2 Ablation Studies
Table 4: Ablation study over the components of TsLLM, as measured by downstream average accuracy on ECG-QA.
Ablation Category Value Accuracy Change (%)
Baseline 59.6
LLM SmolLM2 135M 58.4 -2.0%
LLM SmolLM2 1.7B 60.8 +2.0%
LLM Llama 3.2 3B 62.5 +4.9%
LLM Qwen 2.5 7B 65.0 +9.1%
Adapter MLP 59.3 -0.5%
Adapter Linear 55.8 -6.4%
Encoder VQ-VAE 52.4 -12.1%
We conduct ablation studies on our methodology at a smaller scale, using SmolLM2-360M as the pretrained LLM backbone, with results in Table 4. Performance is measured by accuracy on ECG-QA. These results demonstrate that the size and performance of the backbone LLM matter significantly, suggesting that TsLLM can be scaled up further to improve performance. They also show that simpler adapter layers perform somewhat worse than our cross attention based approach. Finally, using an encoder-decoder pair with quantized latent embeddings currently has a big negative impact on performance, highlighting the importance of precise, high-fidelity encoder and decoder models.
Table 5: Trade-off between reconstruction quality and downstream task performance for different encoder-decoder configurations. Color coding: green indicates improvement, red indicates degradation relative to baseline.
Ablation Baseline Patch size $\beta$ Objective # of layers
Value 16 64 1.0 1.8 AE VQ-VAE 12 48
Recon. MSE 0.217 0.151 0.308 0.173 0.367 0.005 0.432 0.217 0.216
% Change -30% +42% -20% +69% -97% 99% 0% -1%
ECG-QA Acc. 55.9 54.1 53.5 55.5 55.3 54.0 51.8 55.6 54.9
% Change -3.2% -4.3% -0.7% -1.1% -3.4% -7.3% -0.5% -1.8%
Time Series Encoder-Decoder Training
An ablation study on time series encoder-decoder reconstruction performance (Table 5) reveals a critical trade-off between the reconstruction performance of the encoder-decoder pair and downstream performance on practical tasks. We look at the MSE on our VAE validation data across encoder-decoder pairs with different hyperparameter settings, as well as the accuracy of small-scale implementations of TsLLM (SmolLM2-125M, 10B training tokens) trained using each of these pairs. This table makes clear that it is possible to improve the reconstruction ability of the VAE significantly by reducing the patch size or $\beta$, but this leads to more granular representations of the time series that are not as informative or effective when integrated into TsLLM. The representation abstraction can also be increased by increasing $\beta$ or selecting a larger patch size, but this hurts reconstruction, which also hurts performance of TsLLM. We have optimized for a balance between these objectives in order to maximize performance of TsLLM on real-world tasks.
5.3 Performance on Text Generation Tasks
While we have extended the capabilities of the LLM to include time series analysis, our aim is to build a general-purpose model with the same level of intelligence and capabilities as the original pretrained LLM – in this case, Qwen2.5 7B. To this end, we evaluate the performance of TsLLM on standard text generation benchmarks against a fine-tune of Qwen2.5 7B that trains the model on the same text-only datasets we used for training TsLLM. We find nearly identical performance on all benchmarks we tested – specifically MMLU-Pro (53.5 vs 53.5), LiveBench (32.4 vs 32.9), GPQA (36.3 vs 36.4), MATH (68.9 vs 68.1), and GSM8K (89.1 vs 89.2). This indicates that adding time series analysis capabilities to the LLM does not significantly degrade its performance on standard text generation tasks.
6 Discussion
TsLLM augments pretrained LLMs with time series perception, enabling analysis that combines temporal reasoning with natural language understanding. This approach parallels vision-language models, which proved valuable by enabling new capabilities rather than by surpassing specialized computer vision models on traditional benchmarks. Similarly, TsLLM’s contribution lies in enabling capabilities that existing approaches cannot provide: answering questions about temporal patterns, incorporating unstructured context into predictions, and generating natural language explanations.
The contextual forecasting and QA results (Tables 1,2,3) demonstrate that neither frontier LLMs nor traditional time series models can bridge this gap alone. LLMs fail to perceive temporal signals accurately while specialized models cannot leverage textual context, explaining TsLLM’s substantial margins over both. Our ablations further show that performance scales with LLM backbone size, suggesting future base model improvements will directly benefit TsLLM without architectural changes. Notably, optimizing reconstruction fidelity hurts downstream performance; the $\beta$-VAE’s regularized representations outperform a near-perfect autoencoder, indicating the LLM benefits from semantic abstraction over signal-level precision.
Our approach is complementary to specialized time series models rather than competitive with them. TsLLM does not aim to replace established methods for pure numerical prediction tasks, where specialized models remain effective and efficient. Instead, it addresses problems requiring integration of temporal analysis with contextual understanding, domain knowledge, and natural language interaction.
Limitations   Several limitations of TsLLM warrant discussion. First, TsLLM requires substantially more computational resources than traditional time series models, both for training and inference. This is an expected tradeoff when leveraging the capabilities of LLMs. The benefit is that TsLLM can be trained once and applied to diverse tasks through in-context learning, amortizing costs across applications and eliminating the need for task-specific model development. Second, while TsLLM demonstrates competitive performance on traditional benchmarks, it does not achieve state-of-the-art results on pure numerical tasks like forecasting or anomaly detection. Our goal is not to compete with specialized models optimized for specific metrics, but to enable capabilities that those models cannot provide. We argue that the performance gap on traditional benchmarks is acceptable given TsLLM’s unique ability to incorporate contextual information, go beyond the standard time series tasks, and answer questions. Finally, evaluation methodology for time series-language tasks lacks standardization. Unlike pure time series or pure language tasks, there are no established benchmarks for assessing models that bridge both modalities. Developing rigorous, expert-validated benchmarks for tasks like time series question-answering and contextual forecasting should be a priority for the field.
Appendix A Additional Results
A.1 Classification